Suppose the exchange rate is $1.95/£. Let r $ = 7%, r £ = 4%, u = 1.14, d = 0.89, and T = 0.5. Using a 2-step binomial tree, calculate the value of a $2.05-strike European put option on the British pound? Please do NOT answer with Excel. Answer Choices: A. $0.1639 B. $0.1775 C. $0.1745 D. $0.1714 E. $0.1810

EDIT: This question does not need anymore information, everything I have written is all that was provided in the original question.
Answer #1

Suppose the exchange rate is $1.95/£. Let r $ = 7%, r £ = 4%, u = 1.14, d = 0.89, and T = 0.5. Using a 2-step binomial tree, calculate the value of a $2.05-strike European put option on the British pound?

Answer: 2.53 Р 2.22 Р lp А 1.95 1.98 I-P P 1:34 1.54 we have, 6 months. (AST =0-5) European put option with K=2.05 So=1.95 anert_d PS U-d Ya 0-04 , t=0.5/2 20.25 u=1114 • d = d=880 0.89 204 X0:25) 0.89 やん 1-14 -0.89 1.01 -0.89 = 1.14. - 0.84 - 0.48 1

At node e 1.98 0.48 С 0.52 1:54 Day off at 1.980-07 pay off at 1.54 = 2:05 -1.54 -0.51 value atic -0.48x0-07:40-52X0.5) 0-30